OpenStudy (anonymous):

I need help with slopes Find the slope of the line that passes through the pair of points 5,4 9,1

OpenStudy (anonymous):

is there like a easy way to do this problem or do you have to write out a graph and that ?

OpenStudy (anonymous):

use the formula \[m=(y _{2}-y _{1})/(x _{2}-x _{1})\]

OpenStudy (anonymous):

oh yeah hold up

OpenStudy (anonymous):

so 5y 9y 4x 9x

OpenStudy (amistre64):

thats why i dont approve of blindly using a formula; if you dont know how to input the correct data, its useless :)

OpenStudy (anonymous):

well u have to substitute the points in.

OpenStudy (anonymous):

basically it is change in y divided by change in x

OpenStudy (amistre64):

5, 4 9,1 , lets zero out one of the point -5,-4 ------ 0 , 0 now, whatever we do toone point has to be done to the other sooo 5, 4 9, 1 -5,-4 -5, -4 ----- ----- 0 , 0 4 , -3 , slope is just y/x soo; -3/4 is the value we are after

OpenStudy (anonymous):

srry beketso I have to give the point to amistre he explained it better with an steps

OpenStudy (anonymous):

can you also help me on another one but let me first go over this one see if I under stand it

OpenStudy (amistre64):

give it a shot; after a few dozen of these it becomes like second nature

OpenStudy (anonymous):

you take your one y coordinate and subtract it from the other y coordinate .u do the same x coordinate .then u divide the y difference by the x difference

OpenStudy (anonymous):

\[y=3/4 x-9\]

OpenStudy (anonymous):

what about that one ?? and its 3/4 then x

OpenStudy (anonymous):

not 4x

OpenStudy (amistre64):

the form of that equation is called "slope-intercept" since it uses that values of the slope and yintercept to build the equation. y = mx + b is the general setup, such that we can fill in the specifics, m=slope and b = y intercept

OpenStudy (anonymous):

ok so this what it should be layed out to ?

OpenStudy (anonymous):

y= 3/4 -9?

OpenStudy (amistre64):

it should be compared with the general setup y = m x + b y = 3/4 x - 9

OpenStudy (amistre64):

since the m position is for slope, the slope of the line is 3/4

OpenStudy (anonymous):

yes so what do we need to find the slope intercept ?

OpenStudy (anonymous):

the y srry

OpenStudy (anonymous):

well directions say to find the slope and y-intercept

OpenStudy (amistre64):

you want to know how to take the 2 points you gave at the start and form the slope intercept equation, right?

OpenStudy (amistre64):

or, is this a new question?

OpenStudy (anonymous):

this isn't a new ? Im still on the 3/4 ?

OpenStudy (anonymous):

I'm trying to find out what the slope and y-intercept is of y=3/4 x-9

OpenStudy (amistre64):

the form of that equation is called "slope-intercept" since it uses that values of the slope and yintercept to build the equation. y = mx + b is the general setup, such that we can fill in the specifics, m=slope and b = y intercept so, we should be compared with the general setup with the specific setup provided y = m x + b y = 3/4 x - 9 what is the value of m? what is the value of b?

OpenStudy (anonymous):

m = 3/4 and b = -9

OpenStudy (amistre64):

very good so slope is 3/4, and yintercept is -9

OpenStudy (anonymous):

so wouldn't that be your answer ?

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

ok so do you know substitution method ?